OpenStudy (anonymous):

Evaluate the intragural. 3y(7+2y^2)^1/2 dy

OpenStudy (anonymous):

Does anyone have any idea?

OpenStudy (turingtest):

\[u=7+2y^2\]\[du=4ydy\]\[\int3y(7+2y^2)^{1/2}dy=\frac34\int u^{1/2}du\]

OpenStudy (anonymous):

Wait. So how did that 3y in front of the u just dissapear?

OpenStudy (turingtest):

I just pulled it outside the integral sign

OpenStudy (turingtest):

\[\int cf(x)dx=c\int f(x)dx\]

OpenStudy (anonymous):

Even though its 3y(u)^1/2? I thought if it was 3+ you could

OpenStudy (anonymous):

Im just not getting where the y went

OpenStudy (turingtest):

\[u=7+2y^2\]\[du=4ydy\to ydy=\frac14du\]\[\int3y(7+2y^2)^{1/2}dy=3\int (7+2y^2)^{1/2}ydy\]I just pulled the three out and put ydy together, so now I can sub in my expressions for 7+2y^2 and ydy that I got above:\[3\int (7+2y^2)^{1/2}ydy=\frac34\int u^{1/2}du\]let me know if something is still eluding you.

OpenStudy (turingtest):

a little slower on the last step:\[3\int (7+2y^2)^{1/2}ydy=3\int(u)^{1/2}(\frac14du)=\frac34\int u^{1/2}du\]

OpenStudy (anonymous):

I sorta get it I think. Just not sure how ydy becames (1/4du) when du did equal 4ydy. Shouldnt it be (1/4du)y?

OpenStudy (turingtest):

the u sub works such that we need to wind up with \[\int u^ndu\]watch the terms that we get from our substitution\[u=7+2y^2\]\[du=4ydy\]\[ydy=\frac14du\]so we rearrange our integral so that we replace ydy by 1/4du if we had ydu we couldn't integrate, so that would not help

OpenStudy (turingtest):

\[3\int (7+2y^2)^{1/2}(ydy)=3\int(u)^{1/2}(\frac14du)=\frac34\int u^{1/2}du\]see how ydy got replaced?

OpenStudy (anonymous):

Oh I think I get it. Thanks a lot

OpenStudy (turingtest):

u-subs can be tricky at first, but with practice it's a flash :D